# Seemingly unrelated Bayesian additive regression trees for cost-effectiveness analyses in healthcare

### v1

@@ 6 Analysis of the TTCM data
The original dataset had some missing observations — for survey items related to the outcome variables only — which Wiertsema et al. (2019) dealt with through multiple imputation. $17\%$ of patients did not complete any follow-up questionnaires, and hence were missing all information on $q$ and some survey information on $c$ (though hospital records were available for all patients). Additionally, $39\%$ and $7\%$ of respondents were missing some (but not all) survey items related to $c$ and $q$ respectively. As missing data is not the subject of this paper, we will avoid this complication by simply working with one imputed dataset, obtained through predictive mean matching (Vink et al., 2014), and treating that as complete data. Specifically, the imputation is applied to the missing survey items prior to the calculation of $c$ and $q$. It follows that the analysis given here is not directly comparable to the original one, and we do not claim that it is more valid in this regard. We discuss this issue further in Section 7.
We will compare three methods for estimating the treatment effects: (1) suBART, (2) mvBART as implemented in skewBART, and (3) Bayesian linear SUR, with default priors as provided in the Stan user’s guide (Stan Development Team, 2024b). We use $m=100$ trees for the tree-based methods. It is worth noting that we do not impose any restrictions on the sets of covariates associated with each response, for any of these methods. All covariates in Table 4 are used. This means that all trees are allowed to form splitting rules using all covariates for suBART, the single set of multivariate trees are allowed to split on all covariates for mvBART, and all linear regressions for BayesSUR also share all covariates. Following Wiertsema et al. (2019), we do not specify any interaction effects or non-linear terms in the linear predictors for BayesSUR, owing to the difficulty of pre-specifying appropriate functional forms in the presence of a large amount of candidate interactions and the associated challenges in terms of model selection. In any case, Dorie et al. (2019) found that linear models perform poorly in causal settings even when also including interactions, polynomial terms, and regularisation to avoid overfitting. Conversely, BART-based methods are well-equipped to automatically capture low-order interactions and non-linearities (Linero and Yang, 2018; Rocková et al., 2020).
In addition, we evaluate each method again with the set of predictors augmented using propensity scores estimated via probit BART. This procedure is inspired by the univariate ps-BART method proposed by Hahn et al. (2020), which induces a covariate-dependent prior on the regression function and can substantially reduce bias due to regularisation-induced confounding. We first estimate propensity scores for each patient through probit BART, which give the probability that a patient received treatment $1$, conditional on their baseline characteristics, and then add the posterior mean propensity score estimates to the set of predictors $\mathbf{x}_{i}$ used to estimate the conditional expectations $\mathbb{E}[c\>|\>t,\mathbf{x}_{i}]$ and $\mathbb{E}[q\>|\>t,\mathbf{x}_{i}]$ via the chosen model. Thus, we expand the comparison to include what we refer to as ps-suBART, ps-mvBART, and ps-BayesSUR, which are straightforward adaptations of the univariate ps-BART method. Although the use of BART-based propensity scores in conjunction with linear SUR deviates from usual practice in the applied health economics literature, we nonetheless use the same set of propensity scores estimated via probit BART for each method, in order to ensure the comparison is fair in this regard. Expanding the comparison to include versions of each method with and without propensity scores will help to establish the extent to which differences in results are attributable to differences in model specification or due to the inclusion of propensity scores.
6.1 Results of the TTCM data analysis
Figure 7 shows the distribution of the estimated propensity scores. Despite some overlap, it is evident that the treatment groups are quite imbalanced with respect to their baseline characteristics and that some form of covariate adjustment is necessary to avoid biased results. As already described, we reevaluate each model with the estimated propensity scores included as an additional covariate. Following Li et al. (2023), who characterise this as an “approximately Bayesian” procedure, we expect this to lead to results which are more robust to model misspecification. Furthermore, this step makes the models less prone to attributing the effect of confounders to the treatment variable (Hahn et al., 2020).
In Figure 8, we show highest density regions of kernel density estimates of the posterior distributions of $\Delta_{c}$ and $\Delta_{q}$ — obtained through suBART, mvBART, and BayesSUR, as well as their counterpart models which also incorporate the estimated propensity scores as additional predictors — in the form of a cost-effectiveness plane (CEP; see Gabrio et al. (2019) for more details). In general, for both $\Delta_{c}$ and $\Delta_{q}$, the centers of the distributions are shifted further away from 00 and posterior uncertainty is greater when the propensity scores are incorporated. These differences are most pronounced between ps-BayesSUR and BayesSUR and the posterior mean is furthest from the origin under ps-suBART. A notable distinction between suBART and BayesSUR is that the former appears to be more uncertain about $\Delta_{c}$ while the latter appears to be more uncertain about $\Delta_{q}$. Furthermore, the same applies to the comparison between ps-suBART and ps-BayesSUR.
Figure 8: CEPs showing highest density regions of kernel density estimates of the posterior distributions of $\Delta_{c}$ and $\Delta_{q}$ according to each model, with and without propensity scores. The posterior means are indicated by a red dot in each case, the individual draws of $\Delta_{c}$ and $\Delta_{q}$ are shown via grey points, and the contour lines correspond to probability levels of $0.5$, $0.75$, $0.9$, and $0.95$.
In Table 5, we provide summary statistics for $\Delta_{c}$, $\Delta_{q}$, and the INB at a representative value of $\lambda=20000$. We stress that this quantity is not indicative of the cost of the TTCM intervention, which averages €272 per patient (Wiertsema et al., 2019). Rather, this value is indicative of a situation in which decision-makers would be willing to pay €20,000 per additional unit of healthcare-related quality of life, which is much greater as the treatment effect of $\Delta_{q}$ is near-zero. In any case, €20,000 is a commonly used threshold for determining whether an intervention represents value for money in the CEA literature (see e.g., Drummond et al. (2015) and Gabrio et al. (2019). At this chosen $\lambda$ value, we see some considerable differences between ps-suBART and the other methods. This suggests that there may be strong non-linear functional relationships between the covariates and the outcomes (such that suBART benefits from relaxing the linearity assumption of BayesSUR, without requiring pre-specification of the functional forms) and that those relationships may differ for the two outcomes $c$ and $q$ (such that suBART benefits from relaxing the mvBART assumption of a common tree structure). Notably, the estimated treatment effects are markedly smaller in absolute value for each method when propensity scores are excluded and only ps-suBART yields a $95\%$ CI for $\text{INB}_{20000}$ which excludes zero.
Table 5: Posterior means with 95% credible intervals for $\Delta_{c}$, $\Delta_{q}$, and $\text{INB}_{\lambda}$ at a representative value of $\lambda=20000$ for each model, with and without the propensity scores.
Model Mean and 95% CI
$\Delta_{c}$ $\Delta_{q}$ $\text{INB}_{20000}$
ps-suBART $-1371$ $[-2895,189]$ $0.054$ $[-0.020,0.127]$ $2449$ $[210,4635]$
suBART $-880$ $[-2126,345]$ $0.045$ $[-0.019,0.110]$ $1776$ $[-65,3585]$
ps-mvBART $-674$ $[-1859,362]$ $0.042$ $[-0.025,0.112]$ $1516$ $[-282,3415]$
mvBART $-587$ $[-1678,477]$ $0.038$ $[-0.024,0.104]$ $1353$ $[-373,3137]$
ps-BayesSUR $-698$ $[-2166,771]$ $0.093$ $[-0.012,0.203]$ $2560$ $[-228,5395]$
BayesSUR $-191$ $[-1285,916]$ $0.063$ $[-0.016,0.144]$ $1448$ $[-660,3556]$
Rather than relying on a single $\lambda$ value, we also show the probability of cost-effectiveness as a function of the willingness-to-pay $\lambda$ in Figure 9. This plot, called a cost-effectiveness acceptability curve (CEAC; Löthgren and Zethraeus, 2000), is a highly-important tool in guiding the decision of which medical intervention to implement. These probabilities are simply estimated by counting all posterior draws for which $\text{INB}_{\lambda}>0$ and dividing this count by the total number of posterior draws. We see some remarkable differences, depending on which method is used for estimation and whether or not the estimated propensity scores are incorporated. Notably, the estimated probability of cost-effectiveness only exceeds the typical $95\%$ reference level at any $\lambda$ under ps-suBART, suBART, and ps-BayesSUR. Moreover, this probability is consistently larger at all $\lambda$ values for all methods when propensity scores are included. Given that we have reason to believe that ps-suBART, ps-mvBART, and ps-BayesSUR are more accurate than their counterparts, we henceforth discuss only these methods.
A particularly striking aspect of Figure 9 is that ps-suBART is the only method for which the estimated probability of cost-effectiveness is well above $95\%$ for all $\lambda$ values. This threshold is typically regarded as a reasonably high probability of cost-effectiveness, so we would be quite content in asserting that the TTCM is cost-effective, regardless of the value of $\lambda$. In fact, even if decision-makers are unwilling to pay anything (i.e., $\lambda=0$) per unit of effect gained, ps-suBART still reports a probability of $0.97$ of being cost-effective compared to regular care. For ps-BayesSUR, the probabilities are notably lower for small $\lambda$ values, even when including propensity scores. Based on the ps-BayesSUR results, we likely would want to collect more data before committing to a final decision. At higher $\lambda$, the probabilities of cost-effectiveness approach the estimates obtained by ps-suBART. For ps-mvBART, the probabilities follow a similar pattern as the ps-suBART estimates, but remain lower across the full range of $\lambda$ values and never reach the $0.95$ reference level. We are thus lead to a substantially different conclusion, depending on the method used: ps-suBART finds strong evidence for TTCM being cost-effective, while the results from the other two models are less conclusive. As previously alluded to, the results for each method are even less conclusive when the propensity scores are omitted from the set of predictors.
Figure 9: CEACs showing probabilities of cost-effectiveness as a function of $\lambda$ for each model, with and without propensity scores, with a horizontal dotted line at the $0.95$ reference level.
7 Discussion
In this paper, we introduce the suBART model for multivariate outcomes both as means of accounting for non-linearities and interactions in the seemingly unrelated regression framework and as a means of addressing the key limitation of existing multivariate BART approaches which assume a single set of trees, such that the entire response vector is partitioned in the same way by the splitting rules in the ensemble. By modelling each component of the outcome using a univariate BART, suBART captures non-linearities in the relationship between the response and predictors while allowing for and detecting the different subsets of covariates — along with the interactions between them — associated with each response without enforcing common tree structures. We further develop the model to handle multivariate binary outcomes.
The effectiveness of suBART is demonstrated through extensive simulation studies, in which it is shown that the model adequately captures non-linear responses, accurately estimates the covariance structure for multivariate responses, and generally outperforms its main competitors — including other tree-based alternatives and the Bayesian linear SUR — from the points of view of both predictive accuracy and uncertainty calibration. In particular, the extensive evaluation through simulation show that the model exhibits enhanced flexibility compared to its direct multivariate counterpart, mvBART. This flexibility stems from suBART permitting variation in splitting rules across each response, by allowing the trees for each outcome component to differ rather than imposing common tree structures. This enables a more accurate representation, especially when different outcome components depend on distinct sets of predictors.
[…]

### v3

@@ 2.2 Transmural Trauma Care Model (TTCM)
3 The suBART models for continuous and binary responses
3.1 A review of univariate BART
BART was designed to solve the classic regression problem of the form $y_{i}=\mathbb{E}\left[y_{i}\>|\>\mathbf{x}_{i}\right]+\varepsilon_{i}$, where $y_{i}$ is a univariate response variable for observation $i=1,\ldots,n$, $\mathbf{x}_{i}$ is a $p-$dimensional predictor, and $\varepsilon_{i}\sim\text{N}(0,\sigma^{2})$. The idea is to find a flexible approximation for the conditional expectation $\mathbb{E}[y_{i}\>|\>\mathbf{x}_{i}]$ by expressing it as a sum of regression trees. A regression tree consists of two components:
  1. 1.
    A binary tree $\mathcal{T}$, which defines a finite partition $\{\mathcal{A}_{1},\ldots,\mathcal{A}_{h}\}$ of $\mathbb{R}^{p}$ based on the feature space of $\mathbf{X}$, using the available predictors to form splitting rules. In other words, $\mathcal{A}_{1},\ldots,\mathcal{A}_{h}$ are subsets of $\mathbb{R}^{p}$ such that any $\mathbf{x}_{i}\in\mathbb{R}^{p}$ is contained in exactly one $\mathcal{A}_{\ell}$.
  2. 2.
{forest}
$y_{i}=\begin{cases}1&\text{ if }z_{i}>0\\ 0&\text{ otherwise}.\end{cases}$ (3)
As before, we model $z_{i}$ as $z_{i}=\sum^{m}_{t=1}g\left(\mathbf{x}_{i},\mathcal{T}_{t},\mathcal{M}_{t}\right)+\varepsilon_{i}$, with $\varepsilon_{i}\sim\text{N}(0,1)$. This then implies
$\mathbb{E}\left[y_{i}\>|\>\mathbf{x}_{i}\right]=\Pr\left(y_{i}=1\>|\>\mathbf{x}_{i}\right)=\Pr\left(z_{i}>0\>|\>\mathbf{x}_{i}\right)=\Phi\left(\sum_{t=1}^{m}g\left(\mathbf{x}_{i},\mathcal{T}_{t},\mathcal{M}_{t}\right)\right),$
[…]
$=\left[\prod_{t=1}^{m}\prod_{\ell=1}^{h_{t}}\pi\left(\mathcal{\mu}_{t\ell}\>|\>{\mathcal{T}_{t}}\right)\times\pi\left(\mathcal{T}_{t}\right)\right]\times\pi\left(\sigma^{2}\right).$
To achieve conjugacy, it is typically assumed that the residual variance parameter $\sigma^{2}$ follows an inverse-gamma distribution and $\mu_{t\ell}\sim N(0,\sigma_{\mu}^{2})$, with $\sigma_{\mu}^{2}\coloneqq 1/(4\kappa^{2}m)$ being proportional to the number of trees in order to regularise the contribution of each tree. The definition of $\pi(T_{t})$ includes specifying the probability of a non-terminal node as $\alpha(1+\gamma_{t\ell})^{-\beta}$ where $\gamma_{t\ell}$ denotes the depth of that node. The hyperparameters $\alpha$ and $\beta$ take the default values suggested in Chipman et al. (2010) of $0.95$ and $2$, respectively, to favour shallow trees.
[…]
$\pi\left(\mathcal{T}_{t}\>|\>\mathbf{r}_{t},\sigma^{2}\right)$
where new trees and splitting rules are sampled through a Metropolis-Hastings step calculated using the integrated-likelihood for the tree $\mathcal{T}_{t}$ over the leaf parameters $\mathcal{M}_{t}$. See Chipman et al. (2010), Kapelner and Bleich (2016), and Tan and Roy (2019) for further details of the model and additional information the algorithmic implementation on which that of the suBART models is based.
3.2 The suBART model for continuous outcomes
We consider a regression problem of the form
where $\mathbf{y}^{(j)}$ represents the $j$-th component of a $d$-variate outcome, and $(\varepsilon_{i}^{(1)},\ldots,\varepsilon_{i}^{(d)})^{\top}\sim\text{MVN}_{d}(\mathbf{0}_{d},\boldsymbol{\Sigma})$. In principle, different models can be used for each conditional expectation in Equation (4). If the conditional expectations are all assumed to be linear in $\mathbf{x}_{i}$, we obtain the classic SUR model (Zellner, 1962) [^4] A more modern presentation from a Bayesian perspective, closer in spirit to the present work, may be found in Greenberg (2012), Chapter 9. . We instead want to allow the possibility that the conditional expectations are non-linear. Given that the BART model has been shown to be a viable model in the one-dimensional setting, it seems reasonable to expect this viability to extend to the multivariate case. We therefore proceed by assigning an ensemble of regression trees to each $\mathbb{E}[y_{i}^{(j)}\>|\>\mathbf{x}_{i}]$ as follows:
where
  • $\mathcal{T}^{(j)}_{t}$ is a binary tree which defines a finite partition $\{\mathcal{A}^{(j)}_{t\ell}:1\leq\ell\leq h^{(j)}_{t}\}$ of $\mathbb{R}^{p}$. Note that $h^{(j)}_{t}$ is the number of leaf nodes of the tree $\mathcal{T}^{(j)}_{t}$. The collection of all trees pertaining to the $j$-th outcome is denoted by $\boldsymbol{\mathcal{T}}^{(j)}=(\mathcal{T}_{1}^{(j)},\dots,\mathcal{T}_{m}^{(j)})$.
[…]
The model setup is straightforward, and indeed similar to the original BART model in Chipman et al. (2010). Our model is comprised of $d$ univariate BART models, which are linked through correlated error terms $\boldsymbol{\varepsilon}^{(j)}$. Although allowing an outcome-specific number of trees $m_{j}$ is possible, we assume that the number of trees $m$ is common across all $d$ outcomes for simplicity. However, we stress that each outcome is assigned its own separate collection of $m$ trees. Thus, there are $m\times d$ trees in total. This is a key difference compared to other multivariate BART models (McJames et al., 2024; Um et al., 2023), which typically assume that all outcomes share the same $m$ trees and that the leaf parameters associated with each tree are vectors $\boldsymbol{\mu}_{t\ell}$ which follow a $d$-dimensional multivariate normal distribution. Although using $m\times d$ trees in suBART may be computationally prohibitive when $d$ is large, this is rarely the case in causal inference settings, and indeed we have only $d=2$ in the TTCM application.
[…]
We also assume that the leaves of a tree are conditionally independent given the tree structure, i.e.,
$\pi\left(\mathcal{M}^{(j)}_{t}\>|\>\mathcal{T}^{(j)}_{t}\right)=\prod^{h^{(j)}_{t}}_{\ell=1}\pi\left(\mu^{(j)}_{t\ell}\>|\>\mathcal{T}^{(j)}_{j}\right).$
With this setup, prior distributions for $\mathcal{T}^{(j)}_{t},\mu^{(j)}_{t\ell}$, and $\boldsymbol{\Sigma}$ are sufficient to specify the joint prior distribution of all model parameters. The tree structure $\mathcal{T}^{(j)}_{t}$ is assigned the same prior as in the original work by Chipman et al. (2010), with default hyperparameters $\alpha=0.95$ and $\beta=2$ to favour shallow trees and avoid overfitting.
The shrinkage prior used for the leaf node parameters is also aligned with the approach of standard BART. As noted earlier, this prior is formulated conditionally on the tree $\mathcal{T}^{(j)}_{t}$. We assume that each outcome component $\mathbf{y}^{(j)}$ is re-scaled such that $y^{(j)}_{i}\in[-0.5,0.5]$, using its respective observed maximum and minimum values, as outlined by Chipman et al. (2010). This enables us to specify the prior probability that $\mathbb{E}[y^{(j)}_{i}\>|\>\mathbf{x}_{i}]$ lies within the rescaled interval. Consequently, the prior on the leaf node parameters is then
where $\sigma^{(j)}_{\mu}=1/(2\kappa\sqrt{m})$. We suggest $\kappa=2$ as a default choice, which assigns a prior probability of $0.95$ to the event $\{\mathbb{E}[y_{i}^{(j)}\>|\>\mathbf{x}_{i}]\in[-0.5,0.5]\}$. As per the standard BART, having the prior variance of the leaf nodes be inversely proportional to $m$ helps to prevent overfitting. Crucially, the regularisation of each set of $m$ trees per outcome does not depend on the total number of outcomes $d$. Hence, $d$ being large does not incur any additional risk of overfitting.
In the univariate BART model, the error variance is typically assigned an inverse-gamma prior, which is conditionally conjugate and can thus be easily incorporated into a Gibbs sampler. Furthermore, by choosing the hyperparameters accordingly, it is possible to put an informative prior on the error variance. The recommended approach of Chipman et al. (2010) involves obtaining a ‘data-based overestimate’ $\hat{\sigma}^{2}$ of the error variance $\sigma^{2}$ by fitting a linear model via least squares and calculating the standard deviation of the sample residuals [^5] If the number of observations is smaller than the dimension of $\mathbf{X}$, we instead estimate $\hat{\sigma}$ using the LASSO in the glmnet R package (Friedman et al., 2010).. Presumably, the true value of $\sigma^{2}$ is smaller than $\hat{\sigma}^{2}$, since more of the variation in $\mathbf{y}$ can be explained by the flexible BART model than the linear model. Therefore, a large prior probability $\alpha_{\sigma}$ is assigned to the event $\{\sigma^{2}<\hat{\sigma}^{2}\}$ (for example, $\alpha_{\sigma}=0.95$).
We now wish to generalise this idea to the multivariate setting: for all components $j$ of the outcome vector $\left(\mathbf{y}^{(1)},\ldots,\mathbf{y}^{(d)}\right)$, we obtain an overestimate $\hat{\sigma}^{2}_{j}$ of $\sigma^{2}_{j}$ in a similar fashion and want to assign the large prior probability $\alpha_{\sigma}$ to the event $\{\sigma^{2}_{j}<\hat{\sigma}^{2}_{j}\}$. Additionally, we also would like to control the prior on the correlations $\rho_{jk}$, where $j\neq k$. Since there is usually not strong prior information about these correlations, it is preferable for the prior to not be too informative in any direction. Finally, the chosen solution should be computationally tractable and readily incorporated into MCMC samplers. One such candidate is the inverse-Wishart distribution, which is often used as a prior for covariance matrices. As it is a straightforward multivariate generalisation of the inverse-gamma distribution, it facilitates easy computations. However, although it has the desirable property that the prior on $\Sigma$ is invariant to the ordering of the outcome vector $\mathbf{y}_{i}$, the inverse-Wishart prior is unfortunately not well-suited to meet our other aforementioned goals; among other problems, it imposes a strong prior dependency between the variances and the correlations and, when the true variances are small, overestimates variances and induces downward bias towards zero on the correlations (Alvarez et al., 2014).
In order to have more control over the variance parameters, we thus instead adapt an approach by Huang and Wand (2013) and parameterise the covariance matrix $\boldsymbol{\Sigma}$ in a hierarchical fashion as follows:
  • $a_{j}\sim\text{Inv-Gamma}(1/2,1/A^{2}_{j})$, where $A_{j}>0$ is a fixed scale hyperparameter.
  • $\boldsymbol{\Sigma}\>|\>a_{1},\ldots,a_{d}\sim\text{Inv-Wishart}_{d}(\nu+d-1,\mathbf{S}_{0})$, where
    $\mathbf{S}_{0}\coloneqq 2\nu\times\mbox{diag}\left(1/a_{1},\ldots,1/a_{d}\right).$ (7)
As this prior is based on the inverse-Wishart, it shares the property of invariance to the ordering of the outcome vector. A similar prior was proposed by Giannone et al. (2015) in the context of vector autoregressions. Under our prior, the implied prior distribution for the correlations is given by
$\pi\left(\rho_{jk}\right)=\left(1-\rho_{jk}^{2}\right)^{\frac{\nu}{2}-1},\quad\rho_{jk}\in\left(-1,1\right),$ (8)
which matches the prior implied under the standard inverse-Wishart prior (Barnard et al., 2000). Crucially, it does not depend on $A_{1},\ldots,A_{d}$. It is uniform if and only if $\nu=2$. For higher values, the prior increasingly concentrates around zero. Moreover, $\nu=1$ leads to a strongly informative and rarely reasonable prior on the correlations, which favours extreme values near ±$\pm 1$. Thus, we consider $\nu=2$ a reasonable default choice, since we usually do not have any strong prior information on the correlations. It is also worth recalling from Section 2 that the response vector is typically bivariate in the CEA setting which motivated the development of suBART, such that there is only one such correlation parameter.
From the previous definitions, the induced prior for the standard deviations is $\sigma_{j}\sim\text{Half-$t$}(\nu,A_{j})$, which in turn implies a scaled $F$-distribution as the prior for $\sigma_{j}^{2}$. Conversely, an inverse-Wishart prior would imply that the variances have inverse-gamma distributions. As half-$t$ priors on standard deviations are recommended over inverse-gamma priors on variances in univariate settings by Gelman (2006) and others, the construction due to Huang and Wand (2013) is preferable to the inverse-Wishart prior in multivariate settings for similar reasons. Using this construction, we can specify an uninformative prior on each $\sigma_{j}$, by setting a very large value for the associated hyperparameter $A_{j}$. However, in the spirit of the standard BART approach, we prefer to specify informative, data-dependent priors, by choosing each $A_{j}$ in such a way that $\Pr(\sigma_{j}<\hat{\sigma}_{j})=\alpha_{\sigma}$ is ensured a priori.
As well as ensuring that Equation (8) is uniform, another appealing property of specifying $\nu=2$ is the heaviness of the tails of the induced half-$t$ prior on $\sigma_{j}$, which makes the prior more robust to the choice of the scale $A_{j}$ and thereby reduces sensitivity to the data-based overestimate $\hat{\sigma}_{j}$ used for prior calibration. Although Gelman (2006) highlights the special case of the half-Cauchy prior with $\nu=1$ and even heavier tails, we encountered numerical difficulties in our simulation experiments and the TTCM application with $\nu=1$, particularly in situations where the sample size is small, the dimension of $\mathbf{X}$ is large, and the variability of the multivariate outcome is almost entirely explained by $\mathbf{X}$.
We calibrate the prior for each $\sigma_{j}$ by setting up the following equation for each outcome
$\alpha_{\sigma}=\frac{2\Gamma(\frac{\nu+1}{2})}{\Gamma(\frac{\nu}{2})\sqrt{\nu\pi A^{2}_{j}}}\int_{0}^{\hat{\sigma}_{j}}{\left(1+\frac{x^{2}}{\nu A^{2}_{j}}\right)}^{-\frac{\nu+1}{2}}\mathop{}\!\mathrm{d}x$ (9)
and solving it for $A_{j}$, which can be done through numerical root-finding. This expression is the cumulative distribution function of a Half-$t$-distributed random variable with degrees of freedom $\nu$ and scale parameter $A_{j}$, evaluated at $\hat{\sigma}_{j}$. By rewriting Equation (9) in terms of regularised incomplete beta functions, it can be shown that this expression is continuous as well as strictly decreasing in $A_{j}$ and approaches $1$ and 00 as $A_{j}$ approaches 00 and $\infty$, respectively. Thus, the solution for $A_{j}$ exists and is unique.
3.3 Probit suBART
Suppose that we have observed some predictor variables $\mathbf{x}_{i}$ and an associated collection of binary outcome vectors $\mathbf{y}_{i}\in\{0,1\}^{d}$ whose dependence on $\mathbf{x}_{i}$ we want to model. In particular, we do not want to assume that the components of $\mathbf{y}_{i}$ are conditionally independent, given $\mathbf{x}_{i}$; there may be some leftover correlation which is not explained by $\mathbf{x}_{i}$. As in the basic probit BART model, we cast the multivariate version in terms of latent variables $\mathbf{z}_{i}=(z_{i}^{(1)},\ldots,z_{i}^{(d)})$; the construction is exactly as per Equation (3) for each outcome $j$. Then, the probit suBART model is given by
where
The dependence structure of the trees, leaves, and covariance matrix is the same as for the suBART model. The priors for the trees are also exactly the same. Although the other priors are broadly similar, there are some differences that are worth briefly highlighting. The prior for the terminal node parameters $\mu_{t\ell}^{(j)}$ again follows Equation (6), though the calibration of the variance hyperparameter requires more care. Taking ${\sigma^{(j)}_{\mu}}=q_{z}/(\kappa\sqrt{m})$, with $\kappa=2$ as a default choice, we assign a prior probability of $0.95$ to the event $\{\mathbb{E}[z_{i}^{(j)}\>|\>\mathbf{x}_{i}]\in[-q_{z},q_{z}]\}$. On the probability scale, this means that $\{\Pr(y_{i}^{(j)}=1\>|\>\mathbf{x}_{i})\in[\Phi(-q_{z}),\Phi(q_{z})]\}$. For example, when taking $q_{z}=3$ as per Chipman et al. (2010), we assign a prior probability of $0.95$ to the event $\{\Pr(y_{i}^{(j)}=1\>|\>\mathbf{x}_{i})\in[0.0013,0.9987]\}$. This is reasonable for many applications, since extremely small or large probabilities are uncommon.
In the probit setting, $\boldsymbol{\Sigma}$ is a correlation matrix and hence must be positive definite and have all diagonal entries equal to $1$. These restrictions make it difficult to choose a prior for $\boldsymbol{\Sigma}$ which has desirable properties and facilitates easy sampling. Chib and Greenberg (1998) present a prior (and related sampling strategy) which we found to be extremely inefficient in our application. We thus instead adapt an approach by Zhang (2020) (see also Barnard et al. (2000), where most of the following results originate, and Zhang et al. (2006)). We introduce an auxiliary parameter $\mathbf{D}$, which is a $d\times d$ diagonal matrix. We then define $\mathbf{W}\coloneqq\mathbf{D}^{\frac{1}{2}}\boldsymbol{\Sigma}\mathbf{D}^{\frac{1}{2}}$, and assume the prior $\mathbf{W}\sim\text{Inv-Wishart}_{d}(\nu+d-1,\mathbf{I}_{d})$, where $\mathbf{I}_{d}$ is the $d$-dimensional identity matrix. As before, the induced prior on $\boldsymbol{\Sigma}$ is invariant to the ordering of the outcome vector.
where $[\boldsymbol{\Sigma}]_{jj}$ is the $j$-th principle submatrix of $\boldsymbol{\Sigma}$. It can be shown that the marginal prior density for the correlations is again given by Equation (8), as per the suBART model for continuous outcomes above, despite the different priors assumed for $\boldsymbol{\Sigma}$. The hyperparameter $\nu$ plays a similar role as before and we again consider $\nu=2$ to be a reasonable default choice.
4 Posterior inference
This section describes strategies and algorithmic details for conducting posterior inference under suBART and probit suBART for multivariate continuous and multivariate binary outcomes, respectively. For both frameworks, sampling is performed using a Metropolis-within-Gibbs sampler based on their respective priors and model specifications. Given an observed sample $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$, with $\mathbf{y}_{i}\in\mathbb{R}^{d}$, all computations are carried out conditionally on the covariates $\mathbf{X}\in\mathbb{R}^{n\times p}$. Thus, for the sake of readability, we will consider $\mathbf{X}$ fixed and not condition on it explicitly. Additionally, the expression $\pi(\theta\>|\>\boldsymbol{\Theta}\setminus\{\ldots\})$ denotes the conditional distribution of $\theta$ with respect to all parameters except $\theta$ itself and the ones listed within the braces. For example, $\pi(\mathcal{T}^{(j)}_{t}\>|\>\boldsymbol{\Theta}\setminus\{\mathcal{M}^{(j)}_{t}\})$ refers to the distribution of tree $\mathcal{T}^{(j)}_{t}$ given all parameters except $\mathcal{T}^{(j)}_{t}$ and $\mathcal{M}^{(j)}_{t}$. This slightly unusual notation reduces the complexity of the expressions which follow. Given that we routinely condition on all but two parameters, we find it clearer to highlight what is not being conditioned on.
4.1 suBART continuous
For brevity, we define the estimate for $y_{i}^{(j)}$ as $\hat{y}^{(j)}_{i}\coloneqq\linebreak{}\sum_{t=1}^{m}g(\mathbf{x}_{i},\mathcal{T}^{(j)}_{t},\mathcal{M}^{(j)}_{t})$ and $\mathbf{\hat{y}}_{i}=(\hat{y}_{i}^{(1)},\ldots,\hat{y}^{(d)}_{i})^{\top}$, since there are now multiple components of $\mathbf{y}_{i}$. Analogously, we define the residuals from a tree $t$ associated with the $j$-th component as $\mathbf{r}^{(j)}_{t}\coloneqq\{r^{(j)}_{t1},\ldots,r^{(j)}_{tn}\}$, where $r^{(j)}_{ti}\coloneqq y^{(j)}_{i}-\sum_{k\neq t}^{m}g(\mathbf{x}_{i},\mathcal{T}^{(j)}_{k},\mathcal{M}^{(j)}_{k})$. As per the standard BART, we are interested in sampling from the posterior distribution
[…]
which, due to the back-fitting algorithm (Hastie and Tibshirani, 2000) and properties of the multivariate normal distribution, can be obtained from sequential draws from a collection of conditional distributions. In the multivariate continuous outcomes setting, we have that $\mathbf{y}_{i}\sim\text{MVN}_{d}\left(\hat{\mathbf{y}}_{i},\boldsymbol{\Sigma}\right)$. For the following, we will also need the conditional distribution of any component $y^{(j)}_{i}$ given all other components $\mathbf{y}_{i}^{(-j)}$. Using a well-known result (see e.g., Bierens (2004), Section 5.3), this can be found in closed form:
$y^{(j)}_{i}\>|\>\mathbf{y}^{(-j)}_{i},\boldsymbol{\Theta}$ $\sim\text{N}\left(\hat{y}_{i}^{(j)}+\boldsymbol{\Sigma}_{j(-j)}\boldsymbol{\Sigma}_{(-j)(-j)}^{-1}\left(\mathbf{y}^{(-j)}_{i}-\hat{\mathbf{y}}_{i}^{(-j)}\right),\right.$
where $\boldsymbol{\Sigma}_{(-j)(-j)}$ is the submatrix obtained by excluding the $j$-th row and column from $\boldsymbol{\Sigma}$. Analogously, $\boldsymbol{\Sigma}_{j(-j)}$ is the vector obtained by selecting the $j$-th row and excluding the $j$-th column. Using the conditional distribution in Equation (11), Metropolis-Hastings draws from the posterior distribution of trees $\pi(\mathcal{T}^{(j)}_{t}\>|\>\mathbf{r}_{t}^{(j)},\boldsymbol{\Theta}\setminus\{\mathcal{M}_{t}^{(j)}\})$ can be obtained in essentially the same way as for the univariate BART model; see Appendix A of Kapelner and Bleich (2016) for details. Ultimately, as described in Section 3.1, the sampler for the joint posterior distribution of the trees $\mathcal{T}^{(j)}_{t}$ and their parameters $\mathcal{M}_{t}^{(j)}$, for the $j$-th component of $\mathbf{y}^{(1)},\ldots,\mathbf{y}^{(d)}$, is given by successively drawing from $\mathcal{T}^{(j)}_{t}\>|\>\mathbf{r}^{(j)}_{t},\boldsymbol{\Theta}\setminus\{\mathcal{M}_{t}^{(j)}\}$ and $\mathcal{M}^{(j)}_{t}\>|\>\mathbf{r}^{(j)}_{t},\boldsymbol{\Theta}$.
Notably, the algorithm reduces to that of the standard BART approach when $d=1$. Indeed, each draw above can also be viewed as univariate BART — albeit with distinct mean and variance parameters — since it is conditioned on the values of all other components in $\mathbf{Y}$, as illustrated by Equation (11). The full structure of the suBART sampler is given in Algorithm 1, but we first describe the remaining required posterior conditional distributions.
Firstly, the posterior distribution for $\mu^{(j)}_{t\ell}$ is given by
$\mu^{(j)}_{t\ell}\>|\>\boldsymbol{\Theta}\sim\text{N}\left(\left(\frac{{\sigma_{\mu}^{(j)}}^{2}}{v^{(j)}+n_{t\ell}^{(j)}{\sigma_{\mu}^{(j)}}^{2}}\right)\times\left(\sum_{i=1}^{n_{t\ell}^{(j)}}r_{i}^{(j)}-\sum_{i=1}^{n_{t\ell}^{(j)}}u_{i}^{(j)}\right),\frac{v^{(j)}{\sigma_{\mu}^{(j)}}^{2}}{v^{(j)}+n_{t\ell}^{(j)}{\sigma_{\mu}^{(j)}}^{2}}\right),$ (12)
[…]